A nonhomogeneous equation and a particular solution are given. Find a general solution for the equation. 592 0 +20 - 630 =
Answer #1

of Anillary &qh is m² + 2m - 630= 0. os m=-2 4 + 2520 = -2+12524 =-2+2J631 = -1+1631 i lifa C, et(+1+1677 67 +-) ? fe 0 = bes